1: Use 2s complement format to perform the following subtraction 4-5 4 (-5)--1 Answer:
Answer #1

Representation of 4 and -5 in 2's complement
(4)10 = (0100)2
(-5)_{10} = (1011)_2
For subtraction just add them in 2's complement form
(0100)_2 + (1011)_2 = (1111)_2
As MSB in the 2's complement has value -2^n (n = nth bit), -8 in this case.
Converting it to decimal will yield = 1 * -2^3 + 1 * 2^2 + 1*2^1+1*2^0 = -8+4+2+1 = -1
So -1 is the answer.